OpenStudy (firejay5):

8. A record label uses the following function to model the sales of a new release. A(t) = -90t^2 + 8100t. The number of albums sold is a function of time, t, in days. On which day were the most albums sold?

OpenStudy (firejay5):

Is it Day 45???

OpenStudy (anonymous):

so we want the maximum I guess. Take the derivative, setting it equal to 0, and solve for t

OpenStudy (anonymous):

Did you take the derivative?

OpenStudy (anonymous):

Do you know how to take the derivative? Have you been taught that yet? What class is this for? I should be careful as to how I teach you to solve this

OpenStudy (firejay5):

We haven't been taught derivative yet

OpenStudy (anonymous):

ya you wouldn't have been taught that yet. It's so easy that way tho :( \[A(t) = -90t^2 + 8100t\]\[A'(t) = -180t + 8100\]\[0 = -180t + 8100\]\[180t=8100\]\[t=45\] That is the calculus way. I almost want to teach you that. hmm... Algebra 2 method I would graph that on my calculator and use my calculator to find the maximum. Are you allowed to use a graphing calculator for this question?

OpenStudy (firejay5):

Wolframalpha helped me out and was it right

OpenStudy (anonymous):

ya it's right.

OpenStudy (firejay5):

wow! I have 1 more word problem and I will try and figure it out and if not I'll have you help me on it okay

OpenStudy (anonymous):

sure

OpenStudy (firejay5):

thanks